# GOLLuM: Gaussian Process Optimized LLMs – Reframing LLM Finetuning through Bayesian Optimization

### v2

@@ 2.2 Data Representation
BO performance in chemistry is highly sensitive to the choice of data representation, especially given the heterogeneous data types (e.g., categorical reagents, numeric conditions, molecular structures), combinatorial design spaces, and variable numbers of parameters involved – making representation a critical challenge 37. Natural language offers a flexible medium for expressing such optimization problems as textual descriptions, while LLMs can transform these inputs – regardless of type – into unified continuous embeddings. We construct these embeddings through a two-step process:
1. Template Construction: We define each task $t$ as a standardized template: $t=\text{template}(\{\textit{parameters},\textit{values}\})$ where values define the actual conditions of the problem (e.g., reagents used in a chemical synthesis). For single-variable tasks (e.g., molecular optimization), the template reduces to a single textual identifier such as a molecular SMILES string 38, 39. This approach provides a consistent format applicable across a wide range of optimization problems.
@@ 2.3 Gaussian Process with Fixed LLM Embeddings
LLM embeddings can be directly used as input vectors to GPs, which model the output based on observed data. In this setup, the embeddings remain fixed throughout the optimization process – following the approach outlined in BoChemian 31 – and the model’s adaptability comes solely from learning the GP hyperparameters $\theta$. We use a GP prior with a Matérn-5/2 kernel with trainable hyperparameters $\theta=\{\ell,\sigma^{2},\sigma_{n}^{2},c\}$ representing the lengthscale, signal variance, observation noise variance, and constant mean. This approach relies entirely on the pretrained LLM’s embedding space to define input structure – specifically, the relative positioning of points based on their underlying features. The GPs with stationary kernels (such as Matérn-5/2) assume this structure reflects meaningful relationships: points close together in the embedding space are expected to have similar outcomes. However, general pretrained LLMs may not reflect chemical similarities and their representations may not encode the right inductive biases for the task. As a result, the GP can struggle to model the objective effectively in the fixed-feature setting, unless the embedding space already captures relevant patterns. This limitation can be addressed through deep kernel methods, which we describe next.
@@ 2.4 Deep Kernel Gaussian Process
Deep kernel Gaussian processes combine the flexibility of deep neural networks with the principled uncertainty quantification of Gaussian processes. In this approach, the kernel function is composed with a learned feature transformation:
@@ 2.5 LLM-based Deep Kernel
2. PEFT-Adapted LLM: Low-rank adaptation of LLM parameters: $g_{\phi}(\mathbf{x})=\text{LLM}\phi(t)$ where $\phi$ represents the trainable adapter parameters. Parameter efficient finetuning (PEFT) 40, 41, 42 addresses the challenge of adapting large language models by updating a smaller (often several orders of magnitude fewer) number of parameters, typically inserted into or alongside the LLM architecture. We employ Low-Rank Adaptation (LoRA) 41 to preserve potential chemical knowledge captured during pretraining and learn task-specific adaptations, while avoiding catastrophic forgetting or compromising general capabilities.
@@ 2.6 LLM Finetuning as GP Marginal Likelihood Optimization
$\mathcal{L}(\theta,\phi)=\log p(\mathbf{y}|\mathbf{X},\theta,\phi)=-\frac{1}{2}(\mathbf{y}^{\top}\mathbf{K}_{\theta,\phi}^{-1}\mathbf{y}+\log|\mathbf{K}_{\theta,\phi}|+n\log 2\pi)$ (1)
@@ 2.7 Implicit Metric Learning
$\mathcal{L}_{\text{implicit}}(\theta,\phi)\propto\sum_{i,j}w_{ij}\cdot\|g_{\phi}(\mathbf{x}_{i})-g_{\phi}(\mathbf{x}_{j})\|_{2},\quad\begin{cases}\|g_{\phi}(\mathbf{x}_{i})-g_{\phi}(\mathbf{x}_{j})\|_{2}\downarrow\text{ if }\|y_{i}-y_{j}\|\text{ is small}\\ \|g_{\phi}(\mathbf{x}_{i})-g_{\phi}(\mathbf{x}_{j})\|_{2}\uparrow\text{ if }\|y_{i}-y_{j}\|\text{ is large}\end{cases}$ (4)
@@
3 Results & Discussion
3.1 Bayesian Optimization with Fixed LLM Features
Figure 1: BO performance with fixed LLM features as input to GP. Average discovery of high-impact regions of the design space. We show percentage of the top 5% reactions found during the optimization, across all five Buchwald-Hartwig reactions. Domain-specific representations include T5Chem-SMILES43, a pretrained chemistry-related LLM with SMILES input, and DFRP44, a reaction fingerprint.
Building on top of BoChemian31, we first evaluate the effectiveness of LLM embeddings as fixed feature extractors for Bayesian optimization of Buchwald-Hartwig (BH) reactions. Our objective is to optimize the yield of this chemical reaction. The parameters include reaction compounds – 15 reactants, 22 additives, 3 bases, and 4 ligands – totaling a design space of 3955 evaluated reactions, split across five distinct products. We use a variety of publicly available LLMs selected through their base architecture: Encoder-based — ModernBERT 45, UAE 46, MXBAI 47; Encoder–Decoder — Instructor 48, T5 49 and its chemistry-related variant T5Chem 43; Decoder–only — Llama series 50, 51, Mistral series 51, Qwen series 52, 53, and OpenAI embeddings 54.
In Figure 1 we show the performance of all LLM-based representations alongside chemistry-specific baselines: DRFP44 – a reaction fingerprint, and T5Chem-SMILES43 – a domain-specialized LLM leveraging reaction SMILES inputs, aligned with its pretraining. All other representations, including T5Chem, employ templated textual procedures, as described previously.
We use the top 5% coverage metric following 50 BO iterations to evaluate the success of BO in uncovering entire regions of high-valued reactions rather than a single optimum. The motivation is two-fold: (1) identifying a single top-performing reaction is of limited practical value, as the yield difference between the absolute best and other high-performing reactions may be negligible; (2) discovering entire regions of successful reaction conditions enables chemists to select reactions that satisfy additional practical constraints such as cost 55, environmental impact 56 or availability of reagents 57. While the chemistry-specialized baselines excel at discovering high-yield reactions, they require SMILES notation as input, limiting their ability to represent diverse reaction conditions beyond molecular structures.
3.2 What makes a good representation for Bayesian optimization?
Figure 2: Data representations and their success rates in BO. BO performance correlates with GP smoothness, measured as the ratio of learned lengthscale to average pairwise embedding distance.
We observe substantial variation in BO performance across LLM embeddings, prompting analysis of the underlying factors. T5Chem only performs well with inputs resembling its pretraining data (e.g., reaction SMILES), highlighting the limited generality of domain-specialized LLMs 32 and the critical role of representation choice 37. This observation suggests that input representation influences LLM-based BO through two mechanisms: (1) contextual alignment with pretraining helps models better leverage their learned weights, and (2) the resulting embedding structure affects how well the GP can model the objective under a fixed kernel.
To investigate the second effect, we examine how the embedding space structure interacts with the GP’s inductive bias. Specifically, we compute the ratio between the GP’s learned lengthscale and the average pairwise distance between points in the embedding space. This normalized smoothness ratio reflects how far the GP is willing to generalize relative to the scale of the data distribution. As shown in Figure 2 this metric correlates strongly with BO performance ($r=0.92$). A higher ratio indicates that the GP can model the objective with a smoother fit (see Figure 10 in the Appendix), as the embedding space provides a coherent structure that aligns with the kernel’s assumptions. This alignment allows the GP to generalize across broader regions while still resolving performance differences, leading to more effective acquisition decisions. Our analysis complements Papenmeier et al. 58, who argue that longer lengthscales only help when the objective varies slowly enough to be captured by a smooth surrogate. We extend this insight to the representation level, showing that representations inducing such smoothness naturally – without priors or initialization tricks – enable more effective optimization. One might alternatively expect that better GP fit, particularly in high-performing regions, is the primary driver of BO success. However, as shown in Appendix E.3, standard and weighted $R^{2}$ correlate less strongly with performance. While predictive accuracy helps, modeling the objective function and efficiently discovering its optimum are distinct challenges. Our results show that alignment between the representation space and the GP’s inductive bias more directly enables principled optimization.
3.3 Bayesian Optimization with LLM-based Deep Kernels
Figure 3: Comparative analysis of GP-based LLM finetuning. The finetuned models are arranged by the overall performance and relative improvements to their base (fixed embeddings) LLM-GP variants. Chemistry baselines (previous best) included for comparison.
These findings motivate moving beyond fixed embeddings toward joint optimization of both the representation and surrogate model. To address the misalignment between embedding spaces and GP inductive biases, we explore three variants of our LLM-based deep kernel architecture:
1. PLLM: A projection layer applied to fixed LLM embeddings, allowing task-specific transformations with frozen LLM weights.
[…]
In each variant, the deep kernel LLM-GP optimization through shared marginal likelihood enables the representation space to dynamically align with the GP’s assumptions. We now present results supporting this alignment.
GP optimized LLMs as part of the deep kernel GP architecture lead to substantial improvements in performance, increasing the discovery rate of high-performing reactions. Our method consistently outperforms static embeddings across all LLM architectures (Figure 3). $\text{PLLM}\phi$ adaptation of base T5 model now achieves the highest coverage at 42.6%, representing a 74% relative improvement over the fixed-LLM features using the same model (24.45%). Moreover, we observe a 14% increase over the previous best representation (DRFP 37.75%) while not requiring constrained input to reaction SMILES or any domain-specific parametrization. General-purpose LLMs with GP-guided finetuning also surpass chemistry-specialized pretrained LLMs such as T5Chem-SMILES, while using general input through procedural texts. In contrast to previous work 32, these results show that promising BO results can be achieved without requiring domain-specialized models or inputs aligned to their pretraining data. Our model can adapt the input embeddings on-the-fly allowing for increased flexibility. The improvements of GP-finetuned LLMs are also consistent across encoder-only (ModernBERT), encoder-decoder (T5), and decoder-only (Qwen2-7B) architectures. These results validate our earlier analysis on the importance of the alignment between GP’s inductive bias and embedding space structure. In the following section, we analyze how we achieve such alignment.
3.4 Implicit Contrastive Learning and Chemical Interpretability in the Latent Space
Figure 4: Implicit contrastive learning effects with LLM-based deep kernel GPs. A) We observe the progression of the embeddings through pairwise distance histograms of points in separate output classified regions (high yield reactions, low yield reactions, in between). B) Additionally we visualize the new latent space during the optimization procedure in the first, 25th, and last iterations. C) General approach to implicit contrastive learning with GP+LLM architecture. D) Chemical interpretability of the learned latent space. The first three panels show the distribution of reaction outcomes for different reactants (I, Br, Cl aryl halides) across the whole design space (including base-ligand and additive conditions) with colors indicating reaction performance (yield). Following is the projection of the latent space, where we observe reactions clustering based on their aryl halide identity, suggesting that the model captures meaningful chemical relationships. The rightmost section represents the chemical design space of Buchwald-Hartwig reactions used in experiments.
The evolution of the embedding space (Figure 4 A and B) illustrates how the GP’s marginal log-likelihood objective adapts the LLM’s representations during finetuning. Initially, high and low-performing points are mixed in the embedding space. These unstructured representations would typically induce a non-smooth GP fit, as the objective function varies quickly even across nearby points. As optimization progresses, however, the space gradually reorganizes into clearer performance regions (Figure 4 B), achieving a contrastive learning effect through distance-sensitive GP marginal likelihood. The pairwise distance distributions reflect this process – initially overlapping across high-high, high-low, and low-low regions, they gradually separate. This separation is both mathematical and semantical, reflecting the model’s ability to learn chemical relationships that improve optimization (Figure 4 D).
The joint adaptation – with updated LLM embeddings and smoother GP fit, offers greater flexibility than using frozen features. As the structure forms, high-yield reactions become clustered, enabling the GP to guide the acquisition function toward promising regions. Variance estimates remain well-calibrated across the space (visualized with point sizes in Figure 4B), supporting reliable exploration.
Moreover, the learned structure supports interpretability. For instance, in Figure 4D, high-yield iodide-based reactions are consistently separated from lower-yield chloride-based ones. This separation reveals meaningful chemical patterns learned by the model, and highlights how the embedding space captures domain-relevant knowledge that can aid downstream decision-making.
3.5 Additional Benchmarks and Uncertainty Calibration
Figure 5: Benchmarking on various chemistry-related optimization tasks with comparisons to related approaches. A) Visualization of our proposed architectures alongside available previous works (BoChemian 31, LAPEFT 32) and standard GP baseline. B) Optimization results on 19 chemistry-related optimization tasks — 2) reaction optimization (5 Buchwald Hartwig reactions, 4 additive screening reactions, Suzuki Miyaura cross-coupling and catalyst optimization – Bolift 59); 3) general chemistry benchmarks (high-performance liquid chromatography – HPLC setup, oxygen evolution reaction catalysts – OER), vapor diffusion crystallization – Vapdiff) and 4) molecular property optimization. To ensure a fair comparison to LAPEFT 32 we add their best base model (T5Chem with SMILES input) and fix the textual representation to SMILES where possible (BH1-BH5, Additives 1-4, Suzuki-Miyaura, Molecular optimization). However, we note our model’s robustness to a variety of textual input and compare different textual representation of BH reactions (reaction SMILES or textual procedure in B1). For datasets where reaction SMILES are not available (C2 yield, HPLC setup, OER, Vapdiff) we show an example of applied template. C) Predictive ($R^{2}$) and uncertainty estimates (negative log predictive density – NLPD) on BH reactions (60 training points, 20 repeats). Average (across all tasks) optimization performance (top 5% coverage) against benchmarked models.
Finally, we evaluate our model on a variety of optimization tasks in chemistry, from reaction and molecular optimization to the optimization of chemical processes. We provide additional details on the datasets in the Appendix D. All experiments in this and previous sections were run with 10 initial points (randomly selected from the lower median) and 50 BO iterations of batch size 1 (Appendix F).
We compare our results to several related benchmarks: (1) Standard GP with domain-specific representations (DRFP for reactions, molecular fingerprints for property optimization, and vectorized parameters where neither of these are available, as in the case of general chemistry processes optimization tasks). (2) BoChemian 31 which uses fixed LLM embeddings as input to a standard GP. (3) LAPEFT 32 which updates the LLM through supervised finetuning with MSE loss and employs Laplace approximation (LA) for probabilistic modeling. A comprehensive visualization of the different methods is available in Figure 5 A. Having previously shown $\text{PLLM}\phi$ outperforming other variants (PLLM, $\text{LLM}\phi$) we select this approach for all subsequent benchmarking. To ensure a fair comparison to LAPEFT, which states T5Chem with SMILES as the best-performing model, we additionally include this model next to its base variant T5.
Motivated by the insights outlined in 32 we first analyze how the different textual representations influence our model’s performance. For this analysis we compare the results on inputs represented through reaction SMILES or template procedure on BH1-BH5 reactions. We do observe a slight preference of general T5 model to general procedure text and a similar improvement of T5Chem with the structured reaction SMILES compared to a more general input (simplified procedure – Figure 5 B1). However, these differences are still negligible compared to the improvement over fixed features (BoChemian) demonstrating the robustness of our approach to different textual representations (SMILES vs procedure) and LLM pretraining (T5 vs T5Chem). Our method aligns the input through the joint LLM-GP optimization resulting in both adaptive LLM weights – removing the need for domain-pretrained models, and adaptive representations – removing the dependency on the pretraining data format.
[…]
4 Conclusion
This work presents a novel method that reframes LLM finetuning through Bayesian optimization, demonstrating how joint training with Gaussian processes can substantially improve the utility of LLM embeddings for optimization tasks. By leveraging the GP marginal likelihood optimization alongside the representational power of LLMs, we achieve four key benefits: implicit metric learning in the embedding space, principled uncertainty quantification, more effective sampling of promising regions, and seamless adaptability across diverse domains and tasks using readily available LLMs.
Our approach directly addresses two fundamental challenges in BO: (1) designing meaningful representations, which we solve through adaptive LLM embeddings that evolve during optimization, and (2) selecting appropriate kernels, which we handle via deep kernel learning that jointly adapts both the representation space and kernel parameters to the specific task.
By merging LLMs as an integrated part of GPs through deep kernel learning, we enable a structured organization in the embedding space, creating representations that better support sample-efficient optimization. This process occurs without explicit contrastive learning objectives, emerging instead from the GP’s need to model the objective function under uncertainty with trainable LLM input. The consistent improvement across different LLM architectures suggests we have identified a fundamental principle for adapting general pretrained models to specific optimization tasks.
With GP optimized LLMs we offer a principled alternative to prompt-based methods that currently dominate LLM applications. Rather than relying on closed-source model behavior and instruction heuristics that often lack reproducibility, we align the internal representations of LLMs with the requirements of uncertainty-aware decision-making – a critical need in real-world scientific and engineering applications.
Our results across 19 diverse chemical optimization problems demonstrate practical benefits over (1) GP optimization with domain-specialized features (2) static LLM embeddings and (3) decoupled supervised finetuning approaches. These improvements, combined with maintained uncertainty calibration, suggest promising applications beyond chemistry in domains where sample efficiency is crucial and data collection is expensive.
@@ Acknowledgments
We would like to acknowledge Ryan-Rhys Griffiths and Joshua Sin for their invaluable insights, feedback, and constructive discussions throughout this study.
@@
Code and data will be available at https://github.com/schwallergroup/gollum.
@@ Appendix A Related work
Adapting LLMs to specialized tasks through finetuning typically optimizes for predictive accuracy 60, 61. Such domain adaptation neglects the dimension of epistemic uncertainty to indicate when model outputs should not be trusted. In this context, the challenge of extracting reliable uncertainties from LLMs for efficient BO has introduced diverse approaches that fall into four categories.
@@
(1) Prompt-based methods like BOLIFT 59, estimate uncertainty by aggregating multiple LLM responses, while LLAMBO 62 queries LLMs as optimizers. (2) Embedding-based methods, such as BoChemian 31, model GPs on static LLM representations. (3) Surrogate conversion models either predict uncertainty through pretrained in-context regressor 62 or transform LLMs into Bayesian neural networks through PEFT and Laplace approximation (LAPEFT 63 following Bayesian LoRA approach 64). (4) Hybrid approaches like BOPRO 65 combine fixed LLM embeddings for GP-based acquisition optimization with prompt-conditioned LLM generation, using in-context examples to explore the solution space. While promising, these methods face limitations from heavy prompt engineering to post-hoc uncertainty fixes. Our approach offers a solution by integrating uncertainty modeling during training within both the LLM and the surrogate model (GP) through LLM-based deep kernel strategy.
@@
Previous works on integrating LLMs and GPs 37, 59, 31, 32 for chemical optimization primarily use LLMs as fixed feature extractors. Despite LLM’s flexibility in converting diverse parameters into fixed-dimensional representations, these approaches essentially reduce their power to sophisticated encoding tools. The static embeddings struggle to match, let alone surpass, carefully engineered domain-specific features, creating an artificial barrier between LLM expressiveness and GP rigor.
@@
The closest related approach to ours is the one by Kristiadi et al. 32. It incorporates supervised LLM finetuning during optimization (for better prediction), followed by post-hoc Laplace approximation of learned weights (for uncertainty estimates). The sequential approach however, decouples uncertainty quantification from the learning process and optimizes for prediction accuracy (lower MSE loss) rather than optimization itself. Our method fundamentally differs by integrating LLMs directly into the GP framework as a deep kernel, making uncertainty quantification an integral part of the optimization objective through the GP marginal likelihood.
@@
Existing deep kernel methods 66, 67 in chemistry constrain to using graph neural network kernels and domain specific representations. Advancing this concept, we demonstrate that LLM-based kernels provide richer embedding space and easier adaptation to various domains. In the latent space optimization, our approach relates to BO with variational autoencoders (VAEs). For example, 57 use VAE-BO for constrained optimization of molecules while 68 introduce explicit deep metric learning to structure the latent space. We elevate this approach by replacing pretrained VAEs with general purpose LLMs. Moreover, our joint optimization induces implicit metric learning, directly structuring the latent space without the need for explicit contrastive objectives. Importantly, deep kernel learning in low-data regime has long been considered challenging due to issues like overfitting and training instability 69, with some approaches opting for coordinate ascent or bi-level optimization strategies to mitigate these issues 70, 67. In contrast, we build a robust and stable training pipeline using standard backpropagation with separate learning rates for the LLM and GP components, avoiding the need for complex optimization schedules or custom regularization.
@@
Operating in high-dimensional BO (HDBO) spaces (768 features with BERT-based models, 4096 with larger decoder types), our work demonstrates successful optimization that challenges conventional assumptions about HDBO limitations 71. Through our empirical analysis, we identify key principles for effective HDBO: embeddings should structure following the GP’s inductive bias. This organization of the embedding space plays a crucial role in overcoming dimensionality challenges that have traditionally constrained BO approaches. Recent work on HDBO has identified vanishing gradients in GP training and acquisition optimization as a key challenge, and proposed solutions based on informed lengthscale priors and local search heuristics 58. Our approach complements this line of work by focusing on the structure of the representation space itself, showing that smoothness and generalization can emerge naturally when the embedding is well-aligned with the GP kernel. Our observations contribute to emerging approaches in understanding and improving HDBO performance 58, 72.
@@
Within chemistry, high-dimensional features are the default (one-hot encodings 73, fingerprints 74, 75, 76, 44, 77 quantum mechanical descriptors 78, 26) but their selection remains challenging 37. We overcome this problem through unified representation learning as an implicit objective of the optimization strategy itself. As a result, we achieve adaptive features for any optimization at hand, without requiring traditionally used expert-based descriptors.

### v3

@@ 4.2 Data representation
BO performance in chemistry is highly sensitive to the choice of data representation, especially given the heterogeneous data types (e.g., categorical reagents, numeric conditions, molecular structures), combinatorial design spaces, and variable numbers of parameters involved – making representation a critical challenge 22. Natural language offers a flexible medium for expressing such optimization problems as textual descriptions, while LLMs can transform these inputs – regardless of type – into unified continuous embeddings. We construct these embeddings through a two-step process:
1. Template construction: We define each optimization task $t$ as a standardized template: $t=\text{template}(\{\textit{parameters},\textit{values}\})$ where values define the actual conditions of the problem (e.g., reagents used in a chemical synthesis). For single-variable tasks (e.g., molecular optimization), the template reduces to a single textual identifier such as a molecular SMILES string 55, 56. This approach provides a consistent format applicable across a wide range of optimization problems.
@@ 4.3 Gaussian process with fixed LLM embeddings
LLM embeddings can be directly used as input vectors to GPs, which model the output based on observed data. In this setup, the embeddings remain fixed throughout the optimization process – following the approach outlined in BoChemian 42 – and the model’s adaptability comes solely from learning the GP hyperparameters $\theta$. We use a GP prior with a Matérn-5/2 kernel with trainable hyperparameters $\theta=\{\ell,\sigma^{2},\sigma_{n}^{2},c\}$ representing the lengthscale, signal variance, observation noise variance, and constant mean. This approach relies entirely on the pretrained LLM’s embedding space to define input structure – specifically, the relative positioning of points based on their underlying features. The GPs with stationary kernels (such as Matérn-5/2) assume this structure reflects meaningful relationships: points close together in the embedding space are expected to have similar outcomes. However, general pretrained LLMs may not reflect chemical similarities and their representations may not encode the right inductive biases for the task. As a result, the GP can struggle to model the objective effectively in the fixed-feature setting, unless the embedding space already captures relevant patterns. This limitation can be addressed through deep kernel methods, which we describe next.
@@ 4.4 Deep kernel Gaussian process
Deep kernel Gaussian processes 75 combine the flexibility of deep neural networks with the principled uncertainty quantification of Gaussian processes. In this approach, the kernel function is composed with a learned feature transformation:
@@ 4.5 LLM-based deep kernel
2. PEFT-adapted LLM: Low-rank adaptation of LLM parameters: $g_{\phi}(\mathbf{x})=\text{LLM}\phi(t)$ where $\phi$ represents the trainable adapter parameters. Parameter efficient finetuning (PEFT) 76, 77, 78 addresses the challenge of adapting large language models by updating a smaller (often several orders of magnitude fewer) number of parameters, typically inserted into or alongside the LLM architecture. We employ Low-Rank Adaptation (LoRA) 77 to preserve potential chemical knowledge captured during pretraining and learn task-specific adaptations, while avoiding catastrophic forgetting or compromising general capabilities.
@@ 4.6 LLM finetuning as GP marginal likelihood optimization
$\mathcal{L}(\theta,\phi)=\log p(\mathbf{y}|\mathbf{X},\theta,\phi)=-\frac{1}{2}(\mathbf{y}^{\top}\mathbf{K}_{\theta,\phi}^{-1}\mathbf{y}+\log|\mathbf{K}_{\theta,\phi}|+n\log 2\pi).$ (1)
@@ 4.7 Implicit metric learning
[…]
@@
4.8 Joint optimization and computational cost
For deep kernel models, we jointly optimize the GP and LLM parameters by maximizing the marginal likelihood. We use the AdamW optimizer with separate learning rates for the GP hyperparameters (0.2) and the trainable LLM weights (0.002). For larger models, we run experiments on an NVIDIA H100 GPU. LoRA keeps the method lightweight; for T5, fewer than 0.2% of the total model weights are updated during training. We provide full implementation details, including PEFT configurations, learning rate schedules, and hardware specifications, in Supplementary Information section F.3.
@@ Acknowledgments
We would like to acknowledge Joshua Sin and Edvin Fako for their invaluable insights, feedback, and constructive discussions throughout this study.
@@
Code and data available at https://github.com/schwallergroup/gollum.
@@ Appendix A Related work
Adapting LLMs to specialized tasks through finetuning typically optimizes for predictive accuracy 79, 80. Such domain adaptation neglects the dimension of epistemic uncertainty to indicate when model outputs should not be trusted. In this context, the challenge of extracting reliable uncertainties from LLMs for efficient BO has introduced diverse approaches that fall into four categories.
@@
(1) Prompt-based methods like BOLIFT 41, estimate uncertainty by aggregating multiple LLM responses, while LLAMBO and LLM-GO 47, 48 query LLMs as optimizers. (2) Embedding-based methods, such as BoChemian 42, model GPs on static LLM representations. (3) Surrogate conversion models either predict uncertainty through pretrained in-context regressor 47 or transform LLMs into Bayesian neural networks 43 through PEFT and Laplace approximation (LAPEFT 81 following Bayesian LoRA approach 82). (4) Hybrid approaches like BOPRO 44 combine fixed LLM embeddings for GP-based acquisition optimization with prompt-conditioned LLM generation, using in-context examples to explore the solution space. BORA 45 uses a heuristic policy to regulate LLM’s high-level reasoning and its suggestions for new regions of the search space to explore. While promising, these methods face limitations from heavy prompt engineering to post-hoc uncertainty fixes. Our approach offers a solution by integrating uncertainty modeling during training within both the LLM and the surrogate model (GP) through LLM-based deep kernel strategy.
@@
Previous works on integrating LLMs and GPs 22, 41, 42, 43 for chemical optimization primarily use LLMs as fixed feature extractors. Despite LLM’s flexibility in converting diverse parameters into fixed-dimensional representations, these approaches essentially reduce their power to sophisticated encoding tools. The static embeddings struggle to match, let alone surpass, carefully engineered domain-specific features, creating an artificial barrier between LLM expressiveness and GP rigor.
@@
The closest related approach to ours is the one by Kristiadi et al. 43. It incorporates supervised LLM finetuning during optimization (for better prediction), followed by post-hoc Laplace approximation of learned weights (for uncertainty estimates). The sequential approach however, decouples uncertainty quantification from the learning process and optimizes for prediction accuracy (lower MSE loss) rather than optimization itself. Our method fundamentally differs by integrating LLMs directly into the GP framework as a deep kernel, making uncertainty quantification an integral part of the optimization objective through the GP marginal likelihood.
[…]
@@
Existing deep kernel methods 83, 84 in chemistry constrain to using graph neural network kernels and domain specific representations. Advancing this concept, we demonstrate that LLM-based kernels provide richer embedding space and easier adaptation to various domains. In the latent space optimization, our approach relates to BO with variational autoencoders (VAEs). For example, 85 use VAE-BO for constrained optimization of molecules while 86 introduce explicit deep metric learning to structure the latent space. We elevate this approach by replacing pretrained VAEs with general purpose LLMs. Moreover, our joint optimization induces implicit metric learning, directly structuring the latent space without the need for explicit contrastive objectives. Importantly, deep kernel learning in low-data regime has long been considered challenging due to issues like overfitting and training instability 87, with some approaches opting for coordinate ascent or bi-level optimization strategies to mitigate these issues 88, 84. In contrast, we build a robust and stable training pipeline using standard backpropagation with separate learning rates for the LLM and GP components, avoiding the need for complex optimization schedules or custom regularization.
@@
Operating in high-dimensional BO (HDBO) spaces (768 features with BERT-based models, 4096 with larger decoder types), our work demonstrates successful optimization that challenges conventional assumptions about HDBO limitations 89. Through our empirical analysis, we identify key principles for effective HDBO: embeddings should structure following the GP’s inductive bias. This organization of the embedding space plays a crucial role in overcoming dimensionality challenges that have traditionally constrained BO approaches. Recent work on HDBO has identified vanishing gradients in GP training and acquisition optimization as a key challenge, and proposed solutions based on informed lengthscale priors and local search heuristics 90. Our approach complements this line of work by focusing on the structure of the representation space itself, showing that smoothness and generalization can emerge naturally when the embedding is well-aligned with the GP kernel. Our observations contribute to emerging approaches in understanding and improving HDBO performance 90, 91.
@@
Within chemistry, high-dimensional features are the default (one-hot encodings 92, fingerprints 93, 94, 95, 50, 96 quantum mechanical descriptors 49, 20) but their selection remains challenging 22. We overcome this problem through unified representation learning as an implicit objective of the optimization strategy itself. As a result, we achieve adaptive features for any optimization at hand, without requiring traditionally used expert-based descriptors.